Consider a population with values (2,5,7,8)

a. Find the population mean.

b. Find the population variance.

c. Find the population standard deviation.

d. Find the possible samples of size 2 that can be drawn with replacement from this population.


1
Expert's answer
2021-04-15T06:30:25-0400

a. Mean


"\mu=\dfrac{2+5+7+8}{4}=5.5"

b. Variance


"\sigma^2=\dfrac{1}{4}\big((2-5.5)^2+(5-5.5)^2+(7-5.5)^2"

"+(8-5.5)^2\big)=\dfrac{21}{4}=5.25"


c. Standard deviation


"\sigma=\sqrt{\sigma^2}=\sqrt{\dfrac{21}{4}}=\dfrac{\sqrt{21}}{2}\approx2.2913"



d. We have population values "2,5,7,8," population size "N=4" and sample size "n=2." Thus, the number of possible samples which can be drawn without replacement is


"\dbinom{N}{n}=\dbinom{4}{2}=6""\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 Sample & Sample & Sample \ mean \\
 No. & values & (\bar{X}) \\ \hline
 1 & 2,5 & 3.5 \\
 \hdashline
 2 & 2,7 & 4.5 \\
 \hdashline
 3 & 2,8 & 5 \\
 \hdashline
 4 & 5,7 & 6 \\
 \hdashline
 5 & 5,8 & 6.5 \\
 \hdashline
 6 & 7,8 & 7.5 \\
 
 \hline
\end{array}"

The sampling distribution of the sample means.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f & f(\bar{X}) & \bar{X}f(\bar{X})& \bar{X}^2f(\bar{X}) \\ \hline
 7/2 & 1& 1/6 & 7/12 & 49/24 \\
 \hdashline
 9/2 & 1 & 1/6 & 9/12 & 81/24 \\
 \hdashline
 5 & 1 & 1/6 & 10/12 & 100/24 \\
 \hdashline
 6 & 1 & 1/6 & 12/12 & 144/24 \\
 \hdashline
 13/2 & 1& 1/6 & 13/12 & 169/24 \\
 \hdashline
 15/2 & 1 & 1/6 & 15/12 & 225/24 \\
 \hdashline
 Total & 6 & 1 & 66/12 & 768/24 \\ \hline
\end{array}"



"E(\bar{X})=\sum\bar{X}f(\bar{X})=\dfrac{66}{12}=5.5"

The mean of the sampling distribution of the sample means is equal to the

the mean of the population.


"E(\bar{X})=5.5=\mu"


"Var(\bar{X})=\sum\bar{X}^2f(\bar{X})-(\sum\bar{X}f(\bar{X}))^2"

"=\dfrac{768}{24}-(\dfrac{66}{12})^2=\dfrac{42}{24}=\dfrac{7}{4}=1.75"

"\sqrt{Var(\bar{X})}=\sqrt{\dfrac{7}{4}}=\dfrac{\sqrt{7}}{2}\approx1.3229"

Verification:

"Var(\bar{X})=\dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})=\dfrac{\dfrac{21}{4}}{2}(\dfrac{4-2}{4-1})"

"=\dfrac{7}{4}=1.75, True"